# Entanglement Structure of Nonlocal Field Theories

### v1

@@ 2 Entanglement Structure of Nonlocal Fields
To systematically investigate how nonlocality modifies quantum entanglement, we employ a lattice regularization that permits exact computation of entanglement measures. We model a real scalar field as a network of coupled quantum harmonic oscillators on a spatial lattice, with sites labeled by capital Latin indices $M,N$. This discretization serves both as a UV regulator and as an analytically tractable framework for entanglement computation [7, 8, 22]. [^1] For further studies and related computational approaches, see [23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. The system dynamics are governed by the Hamiltonian:
@@ 2.2 Mutual Information and Nonlocal Correlations
Figure 2: Mutual information between two disjoint intervals of equal length ($l_{1}=l_{2}=10$) as a function of the normalized separation $x/L$. Left: For moderate nonlocality parameters ($A=40,60,80$), the mutual information exhibits a gradual decay with separation, highlighting the persistence of correlations at large distances. Right: For strong nonlocality ($A=400,600,800$), the mutual information remains substantial across a wide range of separations, confirming that strong nonlocality drastically slows the decay of quantum correlations.
@@ 3 Holographic Entanglement and the Geometry of Nonlocality
A further redefinition $y=z^{-\omega}$ simplifies the spatial components, yielding a form more amenable to calculating minimal surfaces:
@@
We now consider a boundary subregion $\Omega$ given by a strip of width $l$. The area of a bulk surface $\gamma_{\Omega}$ anchored on $\partial\Omega$ is described by the functional:
@@
Varying this functional yields the equation for the minimal surface. For a 2-dimensional boundary ($d=2$) and a nonlocality exponent $\omega=1$, the problem simplifies due to a conservation law. The minimal surface profile $y(x)$ is then determined by:
@@
where $y_{*}$ is the turning point, the deepest bulk depth reached by the surface. The corresponding area, which gives the holographic entanglement entropy, becomes:
@@
When the probed region is smaller than the nonlocality scale, the connected minimal surface remains close to the boundary ($y_{*}\simeq y_{\infty}$). In this regime, the area scales extensively with the subsystem size, $\text{Area}\propto l$, signaling a volume-law entanglement. This is a direct holographic signature of the highly entangled vacuum state in a nonlocal theory. As shown in Figure 4, the slope of this linear growth increases with $A$, demonstrating that stronger nonlocality enhances entanglement density at short distances.
@@ 4 Conclusion
As a future work, key directions may include extending this analysis to fermionic and gauge-theoretic nonlocal models, exploring dynamical and thermal states, and developing refined holographic frameworks capable of representing the monogamous entanglement and spatial correlation structures that are the hallmark of strongly nonlocal field theories.
@@ Acknowledgments
We gratefully acknowledge M. Alishahiha and M.J. Vasli for their discussions. R. Pirmoradian further thanks Ali Mollabashi, M. Reza Mohammadi Mozafar, and A. Naseh for their valuable early comments and discussions. The work of R. Pirmoradian is based on research funded by the Iran National Science Foundation (INSF) under Project No. 4026389 . Lastly, we recognize the use of AI tools in assisting with text editing and refinement.

### v2

@@ 2 Entanglement Structure of Nonlocal Fields
To systematically investigate how nonlocality modifies quantum entanglement, we employ a lattice regularization that permits exact computation of entanglement measures. We model a real scalar field as a network of coupled quantum harmonic oscillators on a spatial lattice, with sites labeled by capital Latin indices $M,N$. This discretization serves both as a UV regulator and as an analytically tractable framework for entanglement computation [7, 8, 22]. [^1] For further studies and related computational approaches, see [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. The system dynamics are governed by the Hamiltonian:
@@ 2.2 Mutual Information and Nonlocal Correlations
Figure 2: Mutual information between two disjoint intervals of equal length ($l_{1}=l_{2}=10$) as a function of the normalized separation $x/L$ where $L=l_{1}+l_{2}$ is the total length of subsystems. Left: For moderate nonlocality parameters ($A=40,60,80$), the mutual information exhibits a gradual decay with separation, highlighting the persistence of correlations at large distances. Right: For strong nonlocality ($A=400,600,800$), the mutual information remains substantial across a wide range of separations, confirming that strong nonlocality drastically slows the decay of quantum correlations.
@@ 3 Holographic Entanglement and the Geometry of Nonlocality
Using $y=z^{-\omega}$ further simplifies the spatial components, resulting in a form that is more suitable for calculating minimal surfaces:
@@
We consider a boundary subregion $\Omega$ given by a strip of width $l$. The area of a bulk surface $\gamma_{\Omega}$ anchored to $\partial\Omega$ is described by the functional:
@@
Varying this functional yields the equation for the minimal surface, which extends from a deepest point $y_{*}$ to a UV cutoff $y_{\infty}=a^{-\omega}$. This cutoff is the bulk dual to the boundary cutoff in the field theory, giving the range $y_{*}\leqslant y<y_{\infty}$. The corresponding area can be expressed as:
$\text{Area}=2R^{d-2}a^{d-2}A_{0}\int^{y_{\infty}}_{y_{*}}dy\ \dfrac{y^{(2d-3)/\omega}\,y_{*}^{-(d-1)/\omega}}{\sqrt{y^{2(d-1)/\omega}\,\,y_{*}^{-2(d-1)/\omega}-1}}.$ (3.5)
The problem simplifies significantly for a 2-dimensional boundary ($d=2$) with a nonlocality exponent $\omega=1$, due to a conservation law. In this case, the minimal surface profile $y(x)$ is determined by:
@@
where $y_{*}$ is an integration constant identified as the turning point of the surface, defined by $dy/dx_{1}=0$. Consequently, the area simplifies to:
@@
When the size of the probed region is smaller than the nonlocality scale ($l\ll A$), the connected minimal surface remains near the boundary, with its turning point approaching the UV cutoff ($y_{*}\simeq y_{\infty}=a^{-1}$). In this regime, the area exhibits extensive, volume-law scaling with the subsystem size, $\text{Area}\propto l$. This linear dependence is the holographic signature of the highly entangled vacuum state in a nonlocal theory. As shown in Figure 4, the slope of this linear growth increases with the nonlocality scale $A$, indicating that stronger nonlocality enhances the entanglement density at short distances.
@@ 4 Conclusion
Future work should focus on extending this analysis to fermionic and gauge-theoretic nonlocal models, probing dynamical and thermal states, and developing refined holographic frameworks capable of capturing the monogamous entanglement and spatial correlation structures that are characteristic of strongly nonlocal field theories.
Note added
We thank Niko Jokela for valuable correspondence, which highlighted the contrast between our findings, specifically, the absence of a phase transition in the holographic entanglement entropy, and the behavior observed in other holographic nonlocal theories [39, 40]. This discussion helped to clarify the distinctive implications of our geometric setup.
@@ Acknowledgments
We gratefully acknowledge Mohsen Alishahiha and M.Javad Vasli for their discussions. Reza Pirmoradian further thanks Ali Mollabashi, M. Reza Mohammadi Mozafar, and Ali Naseh for their valuable early comments and discussions. The work of Reza Pirmoradian is based on research funded by the Iran National Science Foundation (INSF) under Project No. 4026389. Lastly, we recognize the use of AI tools in assisting with text editing and refinement.